Find the slope and the y-intercept of the line. Write fractional answers in lowest terms. 6y+3x+3=5+3x
The slope is_______________
The y-intercept is ______________

Answer by ewatrrr(24785)


Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
6y+3x+3=5+3x |Solving for y
y = 2/6 = 1/3 OR y = 0*x + 1/3 |slope m is 0, y-intercept is Pt(0,1/3)